OpenStudy (anonymous):

The equation of a circle is x2 + y2 + Cx + Dy + E = 0. If the radius of the circle is decreased without changing the coordinates of the center point, how are the coefficients C, D, and E affected?

OpenStudy (anonymous):

The equation of a circle is x^2 + y^2 + Cx + Dy + E = 0. That is true but there is a much better way of writing that equation: (x-a)^2 + (y-b)^2 = r^2 Try to change x^2 + y^2 + Cx + Dy + E = 0 into something similar to (x-a)^2 + (y-b)^2 = r^2 That is only algebra

OpenStudy (anonymous):

To clear things out (x-a)^2 + (y-b)^2 = r^2 Here (a,b) is the center of the circle and r is the radius

OpenStudy (anonymous):

That was just rude. I am not gonna solve this for you, just guide you so that you can solve it.